I need someones help so bad so it would be a big help if u could help me on this!! ok the problem is Latisha and Maisha are twins. They have a brother who is eleven yeas younger than them and an older sister who is four years older. The sum of the ages of all four siblings is 69. Find Latisha's age. use an algebraic equation
Answer by Paul(988)

2x= sum of the twins
x-11=brother's age
4+x=older sister's age
2x+x-11+4+x=69
4x=76
x=19
Hence, Latisha is 19